Refer to the following information on full-term births in the United States over a given period of time. Type of Birth Number of Births Single birth Twins Triplets Quadruplets 25,900,000 200,000 5000 100 Use this information to estimate the probabilities of the following events. (a) A randomly selected pregnant woman who reaches full term delivers twins. (Give the answer to three significant figures.) (b) A randomly selected pregnant woman who reaches full term delivers quadruplets. (Give the answer to three significant figures.) (c) A randomly selected pregnant woman who reaches full term gives birth to more than a single child. (Give the answer to three significant figures.)
Answer #1

Total births = 25900000 + 200000 + 5000 + 100 = 26105100

a)

P(twins) = n(Twins) / Total births

= 200000 / 26105100

= 0.008

b)

P(quadruplets) = n(quadruplets) / Total births

= 100 / 26105100

= 0.000004

= 0.000 (Rounded to three decimal places)

c)

P(More than a single child) = 1 - P(Single child)

= 1 - n(single child) / Total births

= 1 - [ 25900000 / 26105100]

= 0.008